OpenStudy (anonymous):

The mass of a freight train is 9.98 ⋅ 10^6 kilograms. The mass of an aircraft carrier is 7.3 ⋅ 10^7 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

OpenStudy (jtvatsim):

Welcome to OpenStudy! Have you tried anything so far? Any ideas on where to begin? :)

OpenStudy (anonymous):

I'm Actually Not New xD My Old Account Got Hacked .-. But Thank You Hun cx Umm I was Thing to subtract the decimals and add the Exponents together

OpenStudy (jtvatsim):

Oh that's fun... Well welcome back in that case. :)

OpenStudy (anonymous):

Thank You But Is That Right? to subtract the decimals and add the Exponents together?

OpenStudy (jtvatsim):

Um... I would want to be a bit careful... adding the exponents together would only make sense if we are multiplying. Let's set up the question.

OpenStudy (jtvatsim):

We are trying to find this: (7.3 x 10^7) - (9.96 x 10^6)

OpenStudy (anonymous):

Wait Its Subtraction so i subtract the decimals right?

OpenStudy (jtvatsim):

That's right, but we can only add/subtract the decimals if the power of ten is the same for both.

OpenStudy (anonymous):

Ok so i subract 7300000-996000?

OpenStudy (jtvatsim):

I think you are just missing a zero on each of those, but yes, that's a good idea to convert before subtracting.

OpenStudy (jtvatsim):

73,000,000 - 9,980,000

OpenStudy (anonymous):

Ok so its 63040000 then when i convert it its 6.304 x 10 ^8 ?

OpenStudy (jtvatsim):

You are off by a zero again... :) Let's check how you are counting your zeros... Would you say that 1400 is 1.4 x 10^3 or 1.4 x 10^4 ?

OpenStudy (anonymous):

Huh?

OpenStudy (jtvatsim):

The answer should be 6.304 x 10^7 not 10^8. I want to see how you are deciding what to put for your power of ten.

OpenStudy (jtvatsim):

So as a check, how would you write 1400 in scientific notation (just as a sample question)?

OpenStudy (anonymous):

1.4 x 10 ^3?

OpenStudy (jtvatsim):

Good! I guess you just miscounted the first time for the 6.304 answer. :)

OpenStudy (anonymous):

Yeah xD I have Eye issues to its hard to look at the screen O-o

OpenStudy (jtvatsim):

That probably makes doing scientific calculation miserable... sorry you have to suffer through this part of math. :(

OpenStudy (jtvatsim):

But, yes it looks like you have the correct answer now. When in doubt convert and do normal math. :)

OpenStudy (anonymous):

Its ok :D I just pushes me to work harder but i have a question

OpenStudy (jtvatsim):

Sure, go ahead.

OpenStudy (anonymous):

2^8 ⋅ 5^−5 ⋅ 19^0)^−2 ⋅ 5 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 2^28

OpenStudy (anonymous):

is one of the equations i completed it but i think its wrong :( Would you mind checking it for me?

OpenStudy (jtvatsim):

Whoa... let me try retyping that first to make sure we are doing the same problem.

OpenStudy (anonymous):

This is what i got when i simplified it : = (256 ⋅ 0.00032 ⋅ 1) ⋅ (0.005) ⋅ 2^28 = (149.011611938) ⋅ (0.005) ⋅ 268435456 = (0.74505805969) ⋅ 268435456 = 19.99 with the Vinculum symbol over the 9

OpenStudy (anonymous):

Here is the print shot. its number 2

OpenStudy (jtvatsim):

Oh thank you. that helps. :)

OpenStudy (anonymous):

Your welcome cx

OpenStudy (jtvatsim):

Looks like you might have forget to keep the (0.005) with a power of 4. You have (0.005) should be (0.005)^4 on step one.

OpenStudy (anonymous):

damnit >.< Omg Thank you

OpenStudy (jtvatsim):

lol, your welcome... continuing check...

OpenStudy (anonymous):

i gotta do the whole squation again Dx

OpenStudy (jtvatsim):

lol... looks like there was another error anyways on the computation of 256 * 0.00032. Don't know if that's a comfort or an annoyance. :)

OpenStudy (jtvatsim):

So I agree that step 1 should look like (256 * 0.00032 * 1) * (0.005)^4 * 2^28.

OpenStudy (jtvatsim):

Step 2: (Recalculate first parentheses 256 * ...) (Recalculate 0.005...). 2^28 = 268435456 is correct.

OpenStudy (anonymous):

So wait 149.011611938=(256 * 0.00032 * 1) ?

OpenStudy (jtvatsim):

no, (256* 0.00032 * 1) does not equal 149.011...

OpenStudy (jtvatsim):

Are you doing this with a calculator (I hope)? :)

OpenStudy (anonymous):

Yeah xD i got 0.08192?

OpenStudy (jtvatsim):

Yes, that's better. :)

OpenStudy (anonymous):

Ok xD so part two should look like = (0.08192) * (0.005)^4 * 268435456?

OpenStudy (jtvatsim):

Yep

OpenStudy (anonymous):

the calculator says 6.25e-10 O.o

OpenStudy (jtvatsim):

Yes, this is the calculators way of telling you that the answer is 6.25 x 10^-10, it uses "e" to stand for "x 10^". Weird but true.

OpenStudy (anonymous):

Im confused now

OpenStudy (jtvatsim):

No worries, all you have is (0.08192)*(6.25 x 10^-10)*268435456

OpenStudy (jtvatsim):

6.25 x 10^-10 is 0.000000000625 (nine zeros after decimal point). I'm actually assuming that they wanted you to use scientific notation to solve most of these problems since that makes it easier. Your answers won't be wrong if you convert, but it's easier to make mistakes.

OpenStudy (anonymous):

So it should look like this? = (256 * 0.00032 * 1) * (0.005)^4 * 2^28. = (0.08192) * (0.005)^4 * 268435456 = (0.08192) * (6.25 x 10^-10) * 268435456 = (0.08192) * (0.000000000625) * 268435456

OpenStudy (jtvatsim):

Yes! You've got it.

OpenStudy (anonymous):

Yay cx

OpenStudy (jtvatsim):

Just be careful typing that into a calculator... We'll see if it can handle it.

OpenStudy (anonymous):

i got 5.12e-11=(0.08192) * (0.000000000625)

OpenStudy (jtvatsim):

So did I. :)

OpenStudy (anonymous):

Yay i did it right xD Lol so ten nines are the decimal?

OpenStudy (jtvatsim):

Yes, ten zeros (I knew what you meant) after the decimal.

OpenStudy (anonymous):

xD lol my bad ok

OpenStudy (anonymous):

0.0000000000512?

OpenStudy (jtvatsim):

phew.... lots of zeros, yes looks good.

OpenStudy (anonymous):

Ok xD and then 0.0000000000512 * 268435456?

OpenStudy (jtvatsim):

^Yep, that monster... :)

OpenStudy (anonymous):

xD lol the answer i got was 0.01374389534?

OpenStudy (jtvatsim):

Yep, and it just keeps going... that is the answer.

OpenStudy (jtvatsim):

Now, just invent a robot that can do this automatically...

OpenStudy (anonymous):

xD that is too much work xD

OpenStudy (jtvatsim):

lol, yep

OpenStudy (anonymous):

= (256 * 0.00032 * 1) * (0.005)^4 * 2^28. = (0.08192) * (0.005)^4 * 268435456 = (0.08192) * (6.25 x 10^-10) * 268435456 = (0.08192) * (0.000000000625) * 268435456 = 0.0000000000512 * 268435456 = 0.01374389534 Is the full answer?

OpenStudy (jtvatsim):

Hold on one second... I have to fix something on what we did...

OpenStudy (jtvatsim):

= (256 * 0.00032 * 1)^-2 * (0.005)^4 * 2^28. = (0.08192)^-2 * (0.005)^4 * 268435456 = (0.08192)^-2 * (6.25 x 10^-10) * 268435456 = 149.0116119 * (0.000000000625) * 268435456 = (9.31323 x 10^-8) * 268435456 =

OpenStudy (jtvatsim):

We forgot the -2 exponent on the first parentheses, total oversight on my part, so I recalculated with Excel.

OpenStudy (anonymous):

>.< Lol i knew i was right before xD

OpenStudy (jtvatsim):

Yes, indeed you were... :)

OpenStudy (anonymous):

Im so lost now xD

OpenStudy (jtvatsim):

OK... Let's refresh what happened...

OpenStudy (anonymous):

idk where to pick up from O.o =

OpenStudy (jtvatsim):

= (9.31323 x 10^-8) * 268435456 Let's start here.

OpenStudy (jtvatsim):

Turns out this is 25... wow.

OpenStudy (anonymous):

= (256 * 0.00032 * 1)^-2 * (0.005)^4 * 2^28. = (0.08192)^-2 * (0.005)^4 * 268435456 = (0.08192)^-2 * (6.25 x 10^-10) * 268435456 = 149.0116119 * (0.000000000625) * 268435456 = (9.31323 x 10^-8) * 268435456 = 25?

OpenStudy (jtvatsim):

yes

OpenStudy (anonymous):

Ok

OpenStudy (jtvatsim):

Aw man!... now I see it... We have the right answer, but we should have just left the numbers alone and multiplied the exponents according to the rules. Then, (2^8 x 5^-5 x 19^0)^-2 = 2^-16 x 5^10 (5^-2 / 2^3)^4 = 5^-8 / 2^12 = 5^-8 x 2^-12 and 2^28 is just 2^28. Then we just get 2^-16 x 5^10 x 5^-8 x 2^-12 x 2^28 = 2^-28 x 5^2 x 2^28 = 5^2 = 25

OpenStudy (anonymous):

IM so confused -.-

OpenStudy (anonymous):

@satellite73 Can u help? since you're watching?

OpenStudy (jtvatsim):

nvr mind then, just for future reference, like I said we have the right answer one way or the other. :)

OpenStudy (anonymous):

= (256 * 0.00032 * 1)^-2 * (0.005)^4 * 2^28. = (0.08192)^-2 * (0.005)^4 * 268435456 = (0.08192)^-2 * (6.25 x 10^-10) * 268435456 = 149.0116119 * (0.000000000625) * 268435456 = (9.31323 x 10^-8) * 268435456 = 25? is right thou or no

OpenStudy (jtvatsim):

Yes that is right

OpenStudy (anonymous):

ok xD